the market demand and supply function for milk are: Qd= 58-30.4P and Qs = 16+ 3.2P. If a price floor of $1.75 is implemented, calculate the change in producer surplus. How many surplus units of milk are being produced? If the government purchases all the excess units at $1.75, calculate the milk expenditure by government? Does the increase in producer surplus due to the price floor exceed government spending on excess milk?
1
Expert's answer
2019-03-04T09:58:34-0500

Answer on Question #85720, Economics / Microeconomics


Qd= 58 - 30.4P and Qs = 16 + 3.2P.

In equilibrium Qd = Qs,

58 - 30.4P = 16 + 3.2P,

33.6P = 42,

Pe = $1.25,

Qe = 16 + 3.2*1.25 = 20 units.

Producer surplus before implementing the price floor is:

PS1 = (Qe + Q(0))/2*Pe = (20 +16)/2*1.25 = $22.5.

If a price floor of $1.75 is implemented, the new producer surplus is:

PS2 = (21.6 + 16)/2*1.75 = $32.9.

So, the change in producer surplus is 32.9 - 22.5 = $10.4.

The quantity of surplus units of milk produced is:

Qs(1.75) - Qd(1.75) = 21.6 - 4.8 = 16.8 units.

If the government purchases all the excess units at $1.75, then the milk expenditure by government is:

16.8*1.75 = $29.4.

The increase in producer surplus due to the price floor will not exceed government spending on excess milk.